# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v1

@@ 5.1 Multivariate Regression with $\ell_{2}$-Variational Penalty
To conclude the section would like to underscore the advantage of the combination of using a surrogate loss with exact subgradient following over a dual ascent algorithm tailored for the problem. The dual proximal gradient (DPG) is described and briefly analyzed in Appendix B. To do so, we generated synthetic data of varying levels of variational sparsity as follows. For each sparsity level, let $\mathcal{D}$ denote a distribution whose density function statisfies $\mathcal{D}({\mathbf{x}})\sim\text{e}^{-\lambda\|{\mathbf{x}}\|}$ such that its probability mass within the unit ball is equal to a prescribed fusion probability. We first sample a sequence of Boolean variables $\{\nu_{i}\}$ from the Bernoulli distribution with a predefined probability which designated the chance of fusing two consecutive vectors. When $\nu_{i}=1$, we sampled dual variables $\boldsymbol{\alpha}_{i}$ from $\mathcal{D}$ over unit ball and set ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}=\mathbf{0}$. Otherwise, we sampled ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ from $\mathcal{D}$ outside the unit ball and set $\boldsymbol{\alpha}_{i}\in\partial\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. We then constructed $\mathbf{y}_{i}$ based on the optimality condition, $\mathbf{y}_{i}={\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}+\boldsymbol{\alpha}_{i-1}$. This randomized generation process ensures that ${\mathbf{x}}_{1:n},\boldsymbol{\alpha}_{0:n}$ are primal-dual optimal for $\mathbf{y}_{1:n}$.
@@
We run both DPG and Nariat for 100 iterations each and then merged two consecutive points when the $2$-norm of the variational difference was below a threshold of $10^{-8}$. For each level of variational sparsity we generated $100$ random sequences of vectors with $d=100$ and $n=100$. We measured the $2$-norm between the solution found by the algorithms and the primal-dual optimal solution which was used to generate the synthetic data. The results are shown in Figure 5. The results clearly indicate that Nariat is superior to DPG which was tailored for the problem. The gap between ${\mathbf{x}}^{*}_{1:n}$ and $\hat{{\mathbf{x}}}_{1:n}$ found by Nariat is insensitive of the fusion probability whereas DPG’s performance deteriorates as the level of sparsity grows. Equally if not more important, Nariat exhibits much superior recovery rate of fusion events and in fact seems to slightly improve with the level of sparsity. In contrast, DPG’s rate of recovery rapidly falls to the point that at a sparsity level of $70\%$ almost no fusion events were identified.
@@
Figure 5: Nariat vs. DPG for sparse multivariate discovery using the the $2$-norm .
5.2 Multivariate Regression with $\infty$ norm as variational Penalty
@@
The second vector fusion penalty we consider is the infinity norm of the difference vectors ${\mathbf{x}}_{i}-{\mathbf{x}}_{i+1}$ which entails the problem,
@@
$\min_{{\mathbf{x}}_{1:n},\boldsymbol{\xi}\geq 0}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}.$
We cast the problem as a gradient-based search for the optimum $\xi_{1:n-1}^{*}$. On each gradient step we solve a residual problem in ${\mathbf{x}}$ using the subgradient following procedure with bounded variation from the previous section. Formally, given any $\xi_{1:n-1}\geq 0$ we use the problem given by Eq. 13 to define $\mathcal{L}(\xi_{1:n-1})$ as follows,
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}.$
@@
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\,\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}+\sum_{i=1}^{n-1}\boldsymbol{\alpha}_{i}^{\top}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-\xi_{i}\|\boldsymbol{\alpha}_{i}\|_{1}$
Due to strong duality we can interchange min and max in the minimax problem to get,
@@
$=$ $\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\|_{2}^{2}+\sum_{i=1}^{n-1}\xi_{i}\big{(}1-\|\boldsymbol{\alpha}_{i}\|_{1}\big{)}~{}.$
@@
Therefore $\mathbf{u}\in\partial\mathcal{L}(\xi_{1:n})$, namely, it is a subgradient of $\mathcal{L}$ at $\xi_{1:n-1}$. Again by strong duality, we can obtain the optimal dual variables by first solving the primal problem and then infer the dual variables $\boldsymbol{\alpha}_{:}$ from optimality conditions. Given $\xi_{1:n-1}$ the primal is separable and we thus can apply the subgradient following from 3.5 to obtain both the optimal primal and dual variables. The slack variables $\xi_{1:n-1}$ are updated using gradient stepping. To do so, we derive upper bounds such that $\xi_{i}^{*}\leq D_{i}$ and then apply AdaGrad [5] with an initial per-coordinate learning rate of $D_{i}/\sqrt{2}$. The derivation of the upper bounds is given in Appendinx C.
@@
Figure 6: Nariat vs. DPG for sparse multivariate discovery using the the $\infty$-norm .
@@ 6 High-Order Variational Penalties
Rather than discussing specific forms of discrete difference operators, we analyze a general case stemming from convolution operators. Formally, let $\boldsymbol{\tau}=(-1,\boldsymbol{\tilde{\tau}})$ where $\boldsymbol{\tilde{\tau}}=(\tau_{1},\tau_{2},\ldots,\tau_{k})$ be vectors over $\mathbb{R}^{k+1}$ and $\mathbb{R}^{k}$ respectively. The convolution of ${\mathbf{x}}=(x_{1},\ldots,{\mathbf{x}}_{n})$ with $\boldsymbol{\tau}$ is a vector whose $i$’th entry is, $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=\sum_{m=0}^{k}\tau_{m}\,x_{i-m}~{}.$ Equipped with the aforementioned definitions we now focus on problems of the following general form,
$\min_{\mathbf{x}}\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=k}^{n-1}g_{i}\big{(}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{)}~{}.$ (14)
Note that we can choose $\tau_{0}=-1$ as scaling is a degree of freedom that can be absorbed into $g_{i}$. This constructions implies that $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=(\boldsymbol{\tilde{\tau}}\ast{\mathbf{x}})_{i}-x_{i}$. Here we on purpose shift the indices of $g_{i}$ by $1$ to keep the same notation as used in the algorithms above when plugging in $\boldsymbol{\tilde{\tau}}=(1)~{}.$
We use $\boldsymbol{\tau}\circledast{\mathbf{v}}$ to denote the convolution of a vector ${\mathbf{v}}$ with the reversal of $\boldsymbol{\tau}$ which we define as,
$(\boldsymbol{\tau}\circledast{\mathbf{v}})_{i}=\sum_{m=0}^{k}\tau_{m}v_{i+m}~{}.$
@@
Analogously, we define a sequence of intermediate functions defined as
$f_{i}({\mathbf{x}})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(\boldsymbol{\tau}*{\mathbf{x}})_{j+1}-g_{j}^{*}(\alpha_{j})\right]~{}.$ (15)
This definition means that $f_{i}:\mathbb{R}^{i}\to\mathbb{R}$ and thus formally defined as $f_{i}({\mathbf{x}}_{1:i})$ since ${\mathbf{x}}_{i+1:n}$ is grounded to its optimal vector. As boundary conditions we set $f_{n}({\mathbf{x}})=0$. Similar to the constructions above we introduce an auxiliary function, $\tilde{f}_{i}({\mathbf{x}})=f_{i}({\mathbf{x}})+h_{i}(x_{i})$ which, like $f_{i}$, is a function of ${\mathbf{x}}_{1:i}$
@@
$f_{i}({\mathbf{x}})$ $=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\tilde{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+1}({\mathbf{x}})+h_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$ (16)
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\tilde{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{\tilde{f}_{i+1}(x_{i+1}|\boldsymbol{\mathsf{x}}_{i})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\tilde{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})-\tilde{f}_{i+1}^{*}(\alpha_{i}|\boldsymbol{\mathsf{x}}_{i})~{},$
where $\tilde{f}_{i+1}(\cdot|\boldsymbol{\mathsf{x}}_{i})=\tilde{f}_{i+1}(\boldsymbol{\mathsf{x}}_{i},\cdot)$.
@@
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial\tilde{f}_{i+1}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}~{}.$
@@
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\tilde{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}.$
Notice that the admissible set consists of the full sequence of locally-admissible pairs. Due to strong duality, $S_{i}$ actually defines the point-to-set mapping from a prefix vector $\boldsymbol{\mathsf{x}}_{i}$ to its dual solutions of $f_{i}(\cdot)$. Hence, we can rewrite 15 as,
@@
$=$ $\max_{\boldsymbol{\alpha}_{i:n-1}}\>\sum_{j=i}^{n-1}\left[\alpha_{j}(\boldsymbol{\tau}*\boldsymbol{\mathsf{x}}_{i})_{j+1}-h_{j+1}^{*}(-(\boldsymbol{\tau}\circledast\boldsymbol{\alpha})_{j})-g_{j}^{*}(\alpha_{j})\right]~{}.$
Thus, given any $\boldsymbol{\mathsf{x}}_{i}$ and for any $\boldsymbol{\mathsf{a}}_{i}$ attaining the above maximum we have
$\forall m\in\{0,\ldots,k-1\}:\;(\boldsymbol{\tilde{\tau}}\circledast\boldsymbol{\mathsf{a}}_{i})_{-m}\in\partial_{x_{i-m}}f_{i}(\boldsymbol{\mathsf{x}}_{i})~{}.$
Now, given $S_{i+1}$ we conduct an expansion,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
@@
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\tilde{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(x_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}~{}.$
Using the fact that $\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial\tilde{f}^{*}_{i+1}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})$ we get,
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial\tilde{f}_{i+1}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}.$

### v2

@@ 5.1 Multivariate Regression with $\ell_{2}$-Variational Penalty
To conclude the section we would like to underscore the computational advantage of using a surrogate loss with exact subgradient following over a dual ascent algorithm specifically tailored for the problem. The dual proximal gradient (DPG) is described and shortly analyzed in Appendix B. To do so, we generated synthetic data of varying levels of variational sparsity as follows. For each sparsity level, let $\mathcal{D}$ denote a distribution whose density function satisfies $\mathcal{D}({\mathbf{x}})\sim\text{e}^{-\lambda\|{\mathbf{x}}\|}$ such that its probability mass within the unit ball is equal to a prescribed fusion probability. We first sample a sequence of Boolean variables $\{\nu_{i}\}$ from the Bernoulli distribution with a predefined probability which designated the chance of fusing two consecutive vectors. When $\nu_{i}=1$, we sampled dual variables $\boldsymbol{\alpha}_{i}$ from $\mathcal{D}$ over the unit ball and set ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}=\mathbf{0}$. Otherwise, we sampled ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ from $\mathcal{D}$ outside the unit ball and set $\boldsymbol{\alpha}_{i}\in\partial\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. We then constructed $\mathbf{y}_{i}$ based on the optimality condition, $\mathbf{y}_{i}={\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}+\boldsymbol{\alpha}_{i-1}$. This randomized generation process ensures that ${\mathbf{x}}_{1:n},\boldsymbol{\alpha}_{0:n}$ are primal-dual optimal for $\mathbf{y}_{1:n}$.
@@
We ran both DPG and the Untangled Lariat for 100 iterations each. We then used the resulting solution and merged two consecutive vectors when the $2$-norm of their difference was smaller than $10^{-8}$. For each level of variational sparsity we generated $100$ random sequences of vectors of dimension $d=100$ where each sequence is of length $n=100$. We measured the $2$-norm between the solution found by the algorithms and the primal-dual optimal solution which was used to generate the synthetic data. The results, shown in Figure 6, clearly indicate that Untangled Lariat is superior to DPG which despite the latter being tailored for the problem. The gap between ${\mathbf{x}}^{*}_{1:n}$ and $\hat{{\mathbf{x}}}_{1:n}$ found by Untangled Lariat is insensitive to the fusion probability whereas DPG’s performance deteriorates as the level of sparsity increases. Equally, if not more important, Untangled Lariat exhibits much superior recovery rate of fusion events and in fact seems to slightly improve with the level of sparsity. In contrast, DPG’s rate of recovery drastically falls to the point that at a sparsity level of $70\%$ almost no fusion events were identified.
@@
Figure 6: Untangled Lariat vs. DPG for sparse multivariate discovery with a $2$-norm penalty.
5.2 Multivariate Regression With $\infty$-norm Variational Penalty
@@
The second vector fusion penalty we consider is the infinity norm of the difference vectors ${\mathbf{x}}_{i}-{\mathbf{x}}_{i+1}$ which entails the problem of finding the optimal solution of,
@@
$\min_{{\mathbf{x}}_{1:n},\boldsymbol{\xi}\geq 0}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}~{}.$
We cast the problem as a gradient-based search for the optimum $\xi_{1:n-1}^{*}$.
On each gradient step we solve a residual problem in ${\mathbf{x}}$ using the subgradient following procedure with bounded variation from the previous section. Formally, given any $\xi_{1:n-1}\geq 0$ we use the problem given by Eq. (13) to define $\mathcal{L}(\xi_{1:n-1})$ as follows,
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}~{}.$
@@
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\,\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}+\sum_{i=1}^{n-1}\boldsymbol{\alpha}_{i}^{\top}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-\xi_{i}\|\boldsymbol{\alpha}_{i}\|_{1}~{}~{}.$
From strong duality we can swap the order of min and max in the minimax problem to get,
@@
$=$ $\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\|_{2}^{2}+\sum_{i=1}^{n-1}\xi_{i}\big{(}1-\|\boldsymbol{\alpha}_{i}\|_{1}\big{)}~{}~{}.$
@@
Therefore $\mathbf{u}\in\partial\mathcal{L}(\xi_{1:n})$, namely, it is a subgradient of $\mathcal{L}$ at $\xi_{1:n-1}$. Using strong duality again, we can obtain the optimal dual variables by first solving the primal problem and then infer the dual variables $\boldsymbol{\alpha}_{:}$ from optimality conditions. Given $\xi_{1:n-1}$ the primal is separable and we thus can apply the subgradient following from 3.5 to obtain both the optimal primal and dual variables. The slack variables $\xi_{1:n-1}$ are updated using gradient stepping. To do so, we derive upper bounds such that $\xi_{i}^{*}\leq D_{i}$ and then apply AdaGrad [9] with an initial per-coordinate learning rate of $D_{i}/\sqrt{2}$. The derivation of the upper bounds is given in Appendix C.
@@
Figure 7: Untangled Lariat vs. DPG for sparse multivariate discovery using the the $\infty$-norm .
@@ 6 High-Order Variational Penalties
Rather than discussing specific forms of discrete difference operators, we analyze a general case stemming from convolution operators. Formally, let $\boldsymbol{\tau}=(-1,\boldsymbol{\bar{\tau}})$ where $\boldsymbol{\bar{\tau}}=(\tau_{1},\tau_{2},\ldots,\tau_{k})$ be vectors over $\mathbb{R}^{k+1}$ and $\mathbb{R}^{k}$ respectively. The convolution of ${\mathbf{x}}=(x_{1},\ldots,{\mathbf{x}}_{n})$ with $\boldsymbol{\tau}$ is a vector whose $i$’th entry is, $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=\sum_{m=0}^{k}\tau_{m}\,x_{i-m}$.
Equipped with these definitions we now focus on problems of the following general form,
$\min_{\mathbf{x}}\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=k}^{n-1}g_{i}\big{(}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{)}~{}~{}.$ (14)
Since scaling is a degree of freedom we set $\tau_{0}=-1$ and use the definition above $\boldsymbol{\tau}=(-1,\boldsymbol{\bar{\tau}})$ to get that $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=(\boldsymbol{\bar{\tau}}\ast{\mathbf{x}})_{i}-x_{i}$. Here we on purpose shift the indices of $g_{i}$ by $1$ to keep the same notation as used in the algorithms above when plugging in $\boldsymbol{\bar{\tau}}=(1)$. We use $\boldsymbol{\tau}\circledast{\mathbf{v}}$ to denote the convolution of a vector ${\mathbf{v}}$ with the reversal of $\boldsymbol{\tau}$ which we define as,
$(\boldsymbol{\tau}\circledast{\mathbf{v}})_{i}=\sum_{m=0}^{k}\tau_{m}v_{i+m}~{}~{}.$
@@
Analogously, we define a sequence of intermediate functions,
$f_{i}({\mathbf{x}})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(\boldsymbol{\tau}*{\mathbf{x}})_{j+1}-g_{j}^{*}(\alpha_{j})\right]~{}~{}.$ (15)
This definition means that $f_{i}:\mathbb{R}^{i}\to\mathbb{R}$ and thus formally defined as $f_{i}({\mathbf{x}}_{1:i})$ since ${\mathbf{x}}_{i+1:n}$ is grounded to its optimal vector. As boundary conditions we set $f_{n}({\mathbf{x}})=0$. Similar to the constructions above we introduce an auxiliary function, $f_{i-\frac{1}{2}}({\mathbf{x}})=f_{i}({\mathbf{x}})+h_{i}(x_{i})$ which, like $f_{i}$, is a function of ${\mathbf{x}}_{1:i}$.
@@
$f_{i}({\mathbf{x}})$ $=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\bar{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+1}({\mathbf{x}})+h_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$ (16)
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\bar{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+\frac{1}{2}}(x_{i+1}|\boldsymbol{\mathsf{x}}_{i})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\bar{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})-f_{i+\frac{1}{2}}^{*}(\alpha_{i}|\boldsymbol{\mathsf{x}}_{i})~{}~{},$
where $f_{i+\frac{1}{2}}(\cdot|\boldsymbol{\mathsf{x}}_{i})=f_{i+\frac{1}{2}}(\boldsymbol{\mathsf{x}}_{i},\cdot)$.
@@
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial f_{i+\frac{1}{2}}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}~{}~{}.$
@@
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\bar{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}~{}.$
Note that the admissible set consists of the full sequence of locally-admissible pairs. From strong duality, $S_{i}$ de-facto defines a point-to-set mapping from a prefix vector $\boldsymbol{\mathsf{x}}_{i}$ to its dual solutions of $f_{i}(\cdot)$. Hence, we can rewrite Eq. (15) as,
@@
$=$ $\max_{\boldsymbol{\alpha}_{i:n-1}}\>\sum_{j=i}^{n-1}\left[\alpha_{j}(\boldsymbol{\tau}*\boldsymbol{\mathsf{x}}_{i})_{j+1}-h_{j+1}^{*}(-(\boldsymbol{\tau}\circledast\boldsymbol{\alpha})_{j})-g_{j}^{*}(\alpha_{j})\right]~{}~{}.$
Thus, given $\boldsymbol{\mathsf{x}}_{i}$, for any $\boldsymbol{\mathsf{a}}_{i}$ attaining the above maximum we have
$\forall m\in\{0,\ldots,k-1\}:\;(\boldsymbol{\bar{\tau}}\circledast\boldsymbol{\mathsf{a}}_{i})_{-m}\in\partial_{x_{i-m}}f_{i}(\boldsymbol{\mathsf{x}}_{i})~{}.$
Now, given $S_{i+1}$ we perform the following expansion,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
@@
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\bar{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(x_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}~{}~{}.$
Using the fact that $\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial f^{*}_{i+\frac{1}{2}}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})$ we get,
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial f_{i+\frac{1}{2}}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}~{}.$